OpenStudy (anonymous):

If the hypotenuse of a 45°-45°-90° triangle is 13, what is the length of one of the legs? A. 2004-04-01-04-00_files/i0360004.jpg B. 2004-04-01-04-00_files/i0360003.jpg C. 2004-04-01-04-00_files/i0360001.jpg D. 2004-04-01-04-00_files/i0360002.jpg

OpenStudy (tkhunny):

You really should pull out your Pythagorean Theorem on this one. \(Leg^{2} + Leg^{2} = 13^{2}\)

OpenStudy (butterflydreamer):

heeelllo ^_^ Welcome to Open Study :D Anyway, since your triangle is 45 - 45 - 90, this means it is an ISOSCELES triangle because it has two angles which are equal :) Therefore since we know the hypotenuse (the longest side of the triangle) = 13, we need to find the length of the legs. In order to do so, we need to use Pythagoras' theorem :) a^2 + b^2 = c^2 (where a and b are the legs and c is the hypotenuse) Therefore, let leg = x x^2 + x^2 = 13^2 Now gather like terms and solve for "x" :)

OpenStudy (anonymous):

what will the X's be? 45 and 45?

OpenStudy (radar):

The post states "Therefore, let leg = x"

OpenStudy (radar):

Both legs will be equal so x can represent either one of the legs (or both)

OpenStudy (radar):

Do the math: " Therefore, let leg = x, x^2 + x^2 = 13^2 Now gather like terms and solve for "x" :)"

OpenStudy (radar):

@bcs6300 Welcome to Open Study, a good place to learn.

OpenStudy (radar):

Here is what you need to do to solve this: Step 1. Add the x^2 terms (you have two of them) Step 2. Square the 13 (13 times 13) Step 3. divide both sides by 2. step 4. Take the square root of both sides. Thats it, 4 steps for your answer.